Find the solution of the system of linear equation by using augmented matrix.


\left. \begin{array} { l } { x _ { 1 } - 2 x _ { 2 } = 0 } \\ { 3 x _ { 1 } + 4 x _ { 2 } = - 1 } \\ { 2 x _ { 1 } - x _ { 2 } = 3 } \end{array} \right.

1
Expert's answer
2022-01-17T16:44:30-0500

"x_1-2x_2=0"

"3x_1+4x_2=-1"

"2x_1-x_2=3"


augmented matrix:

"\begin{pmatrix}
 1& -2&|&0 \\
 3& 4&|&-1 \\
2& -1&|&3
\end{pmatrix}"


subtract first row multiplied by 3 from second row; subtract first row multiplied by 2 from third row:

"\begin{pmatrix}
 1& -2&|&0 \\
 0& 10&|&-1 \\
0& 3&|&3
\end{pmatrix}"


so, we get:

"10x_2=-1" and "x_2=1"

this means that the system has not solutions